# Inferring Change Points in High-Dimensional Regression via Approximate Message Passing

### v2

@@ 3.1 State Evolution Characterization of AMP Iterates
Dependence of State Evolution on $\bm{B}$, $\bm{\varepsilon}$, ${\bm{\Psi}}$
The dependence of the state evolution parameters in (10)-(13) on $\bm{B},\bm{\varepsilon}$ can be removed under reasonable assumptions. A standard assumption in the AMP literature (Feng et al., 2022) is:
@@
In Appendix B, we give conditions on $f_{t},g_{t}$, which together with (S0), allow the state evolution equations to be simplified and written in terms of $\bar{{\bm{B}}}\sim{\mathbb{P}}_{\bar{\bm{B}}}$ and $\bar{\varepsilon}\sim{\mathbb{P}}_{\bar{\varepsilon}}$ instead of $(\bm{B},\bm{\varepsilon})$. We believe that the dependence of the state evolution on the signal configuration vector ${\bm{\Psi}}$ is fundamental. Since the entries of ${\bm{\Psi}}$ change value only at a finite number of change points, the state evolution parameters will depend on the limiting fractional values of these change points; see (S1) in Appendix B. This is also consistent with recent change point regression literature, where the limiting distribution of the change point estimators in (Xu et al., 2024) is shown to be a function of the data generating mechanism.
@@ 3.2 Choosing the Denoising Functions $f^{t},g^{t}$
To remove the dependence on ${\bm{\Psi}}$ in the state evolution equations, we can postulate randomness over this variable and take expectations accordingly. Indeed, assume that (S0) holds, and postulate a prior distribution $\pi_{\bar{{\bm{\Psi}}}}$ over $[L]^{n}$. For example, $\pi_{\bar{{\bm{\Psi}}}}$ may be the uniform distribution over all the signal configuration vectors with two change points that are at least $n/10$ apart. We emphasize that our theoretical results do not assume that the true ${\bm{\Psi}}$ is drawn according to $\pi_{\bar{{\bm{\Psi}}}}$. Rather, the prior $\pi_{\bar{{\bm{\Psi}}}}$ allows us to encode any knowledge we may have about the change point locations, and use it to define efficient AMP denoisers $f^{t},g^{t}$. This is done via the following ensemble state evolution recursion, defined in terms of the independent random variables $\bar{{\bm{\Psi}}}\mathbin{\overset{}{\kern 0.0pt\sim}}{\mathbb{P}}_{\bar{{\bm{\Psi}}}}$, $\bar{{\bm{B}}}\mathbin{\overset{}{\kern 0.0pt\sim}}{\mathbb{P}}_{\bar{{\bm{B}}}}$, and $\bar{\varepsilon}\mathbin{\overset{}{\kern 0.0pt\sim}}{\mathbb{P}}_{\bar{\varepsilon}}$.
@@
$\bar{{\bm{\kappa}}}_{\bm{\Theta}}^{t+1,t+1}:=\frac{1}{\delta}\lim_{p\to\infty}\frac{1}{p}\sum_{j=1}^{p}{\mathop{\mathbb{E}}}\left[\left(f_{j}^{t+1}(\bar{\bm{V}}_{\bm{B}}^{t+1})-(\bar{{\bm{\nu}}}^{t+1}_{\bm{\Theta}})^{\top}{\bm{\rho}}^{-1}\bar{\bm{B}}\right)\left(f_{j}^{t+1}(\bar{\bm{V}}_{\bm{B}}^{t+1})-(\bar{{\bm{\nu}}}^{t+1}_{\bm{\Theta}})^{\top}{\bm{\rho}}^{-1}\bar{\bm{B}}\right)^{\top}\right],$ (23)
@@ 3.3 Change Point Estimation and Inference
We now show how the AMP algorithm can be used for estimation and inference of the change points $\{\eta_{1},\dots,\eta_{L^{*}-1}\}$. We first define some notation. Let $\mathcal{X}\subset[L]^{n}$ be the set of all piece-wise constant vectors with respect to $i\in[n]$ with at most $(L-1)$ jumps. This set includes all possible instances of the signal configuration vector ${\bm{\Psi}}$ in $[L]^{n}$, and without loss of generality can account for duplicate signals by increasing $L$. Let the function $U:{\bm{\eta}}\mapsto{\bm{\Psi}}$ denote the one-to-one mapping between change point vectors ${\bm{\eta}}$ and signal configuration vectors ${\bm{\Psi}}$. For a vector $\hat{{\bm{\eta}}}$, we let $|\hat{{\bm{\eta}}}|$ denote its dimension (number of elements).
@@
and $\hat{{\bm{\eta}}}(\bm{\Theta}^{t},\bm{y})=U^{-1}(\hat{{\bm{\Psi}}}(\bm{\Theta}^{t},\bm{y}))$. In (35) we propose a general form for $\hat{{\bm{\Psi}}}$ that is suitable for a generic nonlinear model. A common metric for evaluating the accuracy of change point estimators is the Hausdorff distance (Wang and Samworth, 2018; Xu et al., 2024; Li et al., 2023). The Hausdorff distance between two non-empty subsets $X,Y$ of ${\mathbb{R}}$ is
@@
The proof is given in Appendix D.1. The proof of (31) involves showing that ${d_{H}({\bm{\eta}},\hat{{\bm{\eta}}}(\bm{\Theta}^{t},\bm{y}))}/{n}$ is uniformly pseudo-Lipschitz. For each ${\bm{\eta}}$, Proposition 3 precisely characterizes the asymptotic Hausdorff distance and size errors for a large class of estimators $\hat{{\bm{\eta}}}(\bm{\Theta}^{t},\bm{y})$.
@@
Let $(\bm{\Theta}^{t},\bm{y})$ be as in (LABEL:eq:amp), $(\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},{\bm{\Psi}},\bm{\varepsilon}))$ be as in (7), and ${\bm{\psi}}\in\mathcal{X}$. Suppose the model assumptions in Section 2 as well as $(A1)$$(A3)$ are satisfied. Assume that $(\bm{V},\bm{z})\mapsto p_{\bar{{\bm{\Psi}}}|{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}({\bm{\psi}}|\bm{V},q(\bm{z},{\bm{\Psi}},\bm{\varepsilon}))=:p({\bm{\psi}}|\bm{V},q(\bm{z},{\bm{\Psi}},\bm{\varepsilon}))$ is uniformly pseudo-Lipschitz. Then:
@@ 4 Experiments
In this section, we demonstrate the estimation and inference capabilities of the AMP algorithm in a range of settings, with both synthetic and real data. For synthetic data and for $i\in[n]$, we use i.i.d. Gaussian covariates $\bm{X}_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}(0,\bm{I}_{p}/n)$, and we let $\bm{\varepsilon}_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\mathbb{P}}_{\bar{\varepsilon}}={\sf N}(0,\sigma^{2})$ in the linear and rectified linear models. The denoisers $\{g^{t},f^{t+1}\}_{t\geq 0}$ in the AMP algorithm are chosen according to Proposition 2. The computation of $f^{t}$ is standard (outlined in Appendix C.2 of Arpino et al. (2024b)) and the computation of $g^{t}$ is described in Appendix C.2. The Jacobians of these denoisers are computed using Automatic Differentiation in Python JAX (Bradbury et al., 2018). In all the experiments, we use a uniform prior $\pi_{\bar{\Psi}}$ over all configurations with change points at least $\Delta$ apart, for some $\Delta$ that is a fraction of $n$. Error bars represent one standard deviation, and all experiments are the result of at least $10$ independent trials with $t\leq 15$. Posterior densities are smoothed with a Gaussian kernel of lengthscale $0.5$ for interpretability.
@@ 4.1 Linear Model with Change Points
The two right-most plots in Figure 1 display the approximate posterior over the number of change points, i.e., $\sum_{{\bm{\psi}}\in\mathcal{S}}p_{\bar{{\bm{\Psi}}}|\bar{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}({\bm{\psi}}|\bm{\Theta}^{t},\bm{y})$ where $\mathcal{S}$ contains all configurations with a specified number of change points. We use $p=200,{\mathbb{P}}_{\bar{\bm{B}}}=N({\bm{0}},\bm{I}),\sigma=0.1,\Delta=n/10,L^{*}=3,w=4$, and a uniform prior over the number of change points (zero to three). We observe that the approximate posterior concentrates around the ground truth for moderately large $\delta$.
@@
Figure 2 (left) plots the estimated posterior $p_{\bar{{\bm{\Psi}}}|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|\bm{\Theta}^{t},\bm{y})$ labelled ‘AMP’, and the limiting exact posterior $p_{\bar{{\bm{\Psi}}}|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},{\bm{\Psi}},{\bm{\varepsilon}}))$ labelled ‘Theory’, averaged over $30$ trials with true change point locations at $n/3$ and $158n/$ 15. The posterior is computed over the grid $[0,1]$ subsampled by a factor of ten. The experiment uses $p=800,{\mathbb{P}}_{\bar{\bm{B}}}={\sf N}({\bm{0}},\bm{I}),\sigma=0.1,\Delta=n/5,L=L^{*}=3$, with a uniform prior over all valid configurations. We observe that the estimated posterior closely matches the limiting posterior and concentrates around the ground truth as $\delta$ increases.
@@
Figures 3(a) and 3(b) compare the performance of AMP against four state-of-the-art algorithms: the dynamic programming ($\mathtt{DP}$) approach in (Rinaldo et al., 2021); dynamic programming with dynamic updates $(\mathtt{DPDU})$ (Xu et al., 2024); divide and conquer dynamic programming ($\mathtt{DCDP}$) (Li et al., 2023); and the complementary-sketching-based algorithm $\mathtt{charcoal}$ from Gao and Wang (2022). Hyperparameters are chosen using cross validation (CV), as outlined in Section A.1 of Li et al. (2023). The first three algorithms, designed for sparse signals, combine LASSO-type estimators with partitioning techniques based on dynamic programming. The $\mathtt{charcoal}$ algorithm is designed for the setting where the difference ${\bm{\beta}}^{(\eta_{\ell})}-{\bm{\beta}}^{(\eta_{\ell+1})}$ between adjacent signals is sparse. None of these algorithms uses a prior on the change point locations, unlike AMP which can flexibly incorporate both priors via ${\mathbb{P}}_{\bar{\bm{B}}}$ and $\pi_{\bar{{\bm{\Psi}}}}$. Figure 3(a) uses $p=200,\sigma=0.1,\Delta=n/10,L^{*}=L=3$ and a sparse Bernoulli-Gaussian signal prior ${\mathbb{P}}_{\bar{\bm{B}}}=0.5{\sf N}({\bm{0}},\delta\bm{I})+0.5\delta_{\bm{0}}$. AMP assumes no knowledge of the true sparsity level 0.5 and estimates the sparsity level using CV over a set of values not including the ground truth (details in Appendix E). Figure 3(b) uses $p=300,\Delta=n/10,L^{*}=L=3$ and a Gaussian sparse difference prior with sparsity level $0.5$ (described in (120)). AMP is run assuming a mismatched sparsity level of 0.9 and a mismatched magnitude for the sparse difference vector (details in Appendix E). Figures 3(a) and 3(b) show that AMP consistently achieves the lowest Hausdorff distance among all algorithms and outperforms most algorithms in runtime. Figures 8 and 9 in Appendix E show results from an additional set of experiments comparing AMP with $\mathtt{DCDP}$ (the fastest algorithm in Figure 3(a)). Figure 8 shows the performance of AMP with different change point priors $\pi_{\bar{{\bm{\Psi}}}}$ and suboptimal denoising functions, e.g., soft thresholding. Figure 9 demonstrates the favourable runtime scaling of AMP over $\mathtt{DCDP}$ with respect to $p$, due to the LASSO computations involved in $\mathtt{DCDP}$.
@@
In Figure 4, we consider noiseless compressed sensing, where the signals $\{{\bm{\beta}}^{(i)}\}_{i\in[n]}$ are rotated versions of a $(255,255)$ sparse grayscale image used by Schniter and Rangan (2014). The fraction of nonzero components in the image is 8645/50625. We downsample the original image by a factor of three and flatten, yielding an operation dimension of $p=85^{2}=7225$. We set $\{{\bm{\beta}}^{(i)}\}_{i=1}^{0.3n}$ to be the image, $\{{\bm{\beta}}^{(i)}\}_{i=0.3n}^{0.7n}$ to be a $30^{\circ}$ rotated version, and $\{{\bm{\beta}}^{(i)}\}_{i=0.7n}^{n}$ to be a $45^{\circ}$ rotated version. We run AMP with a Bernoulli-Gaussian prior ${\mathbb{P}}_{\bar{\bm{B}}}$, $\Delta=n/4$ and $L=L^{*}=3$. Figure 4 shows image reconstructions along with the approximate posterior $p_{\bar{{\bm{\Psi}}}|\bar{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|\bm{\Theta}^{t},\bm{y})$. The approximate posterior concentrates around the true change point locations as $\delta$ increases, even when the image reconstructions are approximate. The experiment took one hour to complete on an Apple M1 Max chip, whereas competing algorithms did not return an output within 2.5 hours, due to the larger signal dimension compared to Figure 3(a) ($p=7225$ vs $p=200$).

### v3

@@ 3.1 State Evolution Characterization of AMP Iterates
Parametric Dependence of State Evolution on $\bm{B}$, $\bm{\varepsilon}$, $\bm{\Psi}$}
The parametric dependence of the state evolution parameters in (10)-(13) on $\bm{B},\bm{\varepsilon}$ can be removed under reasonable assumptions. A standard assumption in the AMP literature (Feng et al., 2022) is:
@@
In Appendix B, we give conditions on $f_{t},g_{t}$, which together with (S0), allow the state evolution equations to be simplified and written in terms of $\bar{{\bm{B}}}\sim{\mathbb{P}}_{\bar{\bm{B}}}$ and $\bar{\varepsilon}\sim{\mathbb{P}}_{\bar{\varepsilon}}$ instead of $(\bm{B},\bm{\varepsilon})$. We believe that the parametric dependence of the state evolution matrices on the signal configuration vector $\bm{\Psi}$} is fundamental. Since the entries of $\bm{\Psi}$} change value only at a finite number of change points, the state evolution parameters will depend on the limiting fractional values of these change points; see (S1) in Appendix B. This is also consistent with recent change point regression literature, where the limiting distribution of the change point estimators in (Xu et al., 2024) is shown to be a function of the data generating mechanism.
@@ 3.2 Choosing the Denoising Functions $f^{t},g^{t}$
To remove the parametric dependence on $\bm{\Psi}$} in the state evolution equations, we can postulate randomness over this variable and take expectations accordingly. Indeed, assume that (S0) holds, and postulate a prior distribution $\pi_{\bar{{\bm{\Psi}}}}$ over $L]^{n}$. For example, $\pi_{\bar{{\bm{\Psi}}}}$ may be the uniform distribution over all the signal configuration vectors with two change points that are at least $n/10$ apart. We emphasize that our theoretical results do not assume that the true $\bm{\Psi}$} is drawn according to $\pi_{\bar{{\bm{\Psi}}}}$. Rather, the prior $\pi_{\bar{{\bm{\Psi}}}}$ allows us to encode any knowledge we may have about the change point locations, and use it to define efficient AMP denoisers $f^{t},g^{t}$. This is done via the following ensemble state evolution recursion, defined in terms of the independent random variables $\bar{{\bm{\Psi}}}\mathbin{\overset{}{\kern 0.0pt\sim}}{\mathbb{P}}_{\bar{{\bm{\Psi}}}}$, $\bar{{\bm{B}}}\mathbin{\overset{}{\kern 0.0pt\sim}}{\mathbb{P}}_{\bar{{\bm{B}}}}$, and $\bar{\varepsilon}\mathbin{\overset{}{\kern 0.0pt\sim}}{\mathbb{P}}_{\bar{\varepsilon}}$.
@@
$\bar{{\bm{\kappa}}}_{\bm{\Theta}}^{t+1,t+1}:=\frac{1}{\delta}\lim_{p\to\infty}\frac{1}{p}\sum_{j=1}^{p}{\mathop{\mathbb{E}}}\left[\left(f_{j}^{t+1}(\bar{\bm{V}}_{\bm{B}}^{t+1})-(\bar{{\bm{\nu}}}^{t+1}_{\bm{\Theta}})^{\top}{\bm{\rho}}^{-1}\bar{\bm{B}}\right)\right$.
$\hskip 199.16928pt\left.\left(f_{j}^{t+1}(\bar{\bm{V}}_{\bm{B}}^{t+1})-(\bar{{\bm{\nu}}}^{t+1}_{\bm{\Theta}})^{\top}{\bm{\rho}}^{-1}\bar{\bm{B}}\right)^{\top}\right],$ (23)
@@ 3.3 Change Point Estimation and Inference
[…]
@@
and $\hat{{\bm{\eta}}}(\bm{\Theta}^{t},\bm{y})=U^{-1}(\hat{{\bm{\Psi}}}(\bm{\Theta}^{t},\bm{y}))$. In (35) we propose a general form for $\hat{{\bm{\Psi}}}$ that is suitable for a generic nonlinear model. A common metric for evaluating the accuracy of change point estimators is the Hausdorff distance (Wang and Samworth, 2017; Xu et al., 2024; Li et al., 2023). The Hausdorff distance between two non-empty subsets $X,Y$ of $\mathbb{R}$} is
@@
The proof is given in Appendix D.1. To prove (31), we show that ${d_{H}({\bm{\eta}},\hat{{\bm{\eta}}}(\bm{\Theta}^{t},\bm{y}))}/{n}$ is uniformly pseudo-Lipschitz. For each $\bm{\eta}$}, Proposition 3 precisely characterizes the asymptotic Hausdorff distance and size errors for a large class of estimators $\hat{{\bm{\eta}}}(\bm{\Theta}^{t},\bm{y})$.
@@
Let $(\bm{\Theta}^{t},\bm{y})$ be as in (LABEL:eq:amp), $(\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},{\bm{\Psi}},\bm{\varepsilon}))$ be as in (7), and ${\bm{\psi}}\in\mathcal{X}$. Suppose the model assumptions in Section 2 as well as (A1)(A1)$(A3)$ are satisfied. Assume that $(\bm{V},\bm{z})\mapsto p_{\bar{{\bm{\Psi}}}|{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}({\bm{\psi}}|\bm{V},q(\bm{z},{\bm{\Psi}},\bm{\varepsilon}))=:p({\bm{\psi}}|\bm{V},q(\bm{z},{\bm{\Psi}},\bm{\varepsilon}))$ is uniformly pseudo-Lipschitz. Then:
@@ 4 Experiments
In this section, we demonstrate the estimation and inference capabilities of the AMP algorithm in a range of settings, with both synthetic and real data. For synthetic data, for $i\in[n]$, we use i.i.d. Gaussian covariates $\bm{X}_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}(0,\bm{I}_{p}/n)$, and we let $\bm{\varepsilon}_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\mathbb{P}}_{\bar{\varepsilon}}={\sf N}(0,\sigma^{2})$ in the linear and rectified linear models. The denoisers $\{g^{t},f^{t+1}\}_{t\geq 0}$ in the AMP algorithm are chosen according to Proposition 2. The computation of $f^{t}$ is standard (outlined in Appendix C.2 of Arpino et al. (2024b)) and the computation of $g^{t}$ is described in Appendix C.2. The Jacobians of these denoisers are computed using Automatic Differentiation in Python JAX (Bradbury et al., 2018). In all the experiments, we use a uniform prior $\pi_{\bar{\Psi}}$ over all configurations with change points at least $\Delta$ apart, for some $\Delta$ that is a fraction of $n$. Error bars represent one standard deviation, and all experiments are the result of at least $10$ independent trials with $t\leq 15$. Posterior densities are smoothed with a Gaussian kernel of lengthscale $0.5$ for interpretability.
@@ 4.1 Linear Model with Change Points
The two right-most plots in Figure 1 display the approximate posterior over the number of change points, that is, $\sum_{{\bm{\psi}}\in\mathcal{S}}p_{\bar{{\bm{\Psi}}}|\bar{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}({\bm{\psi}}|\bm{\Theta}^{t},\bm{y})$ where $\mathcal{S}$ contains all configurations with a specified number of change points. We use $p=200,{\mathbb{P}}_{\bar{\bm{B}}}=N({\bm{0}},\bm{I}),\sigma=0.1,\Delta=n/10,L^{*}=3,w=4$, and a uniform prior over the number of change points (zero to three). We observe that the approximate posterior concentrates around the ground truth for moderately large $\delta$.
@@
Figure 2 (left) plots the estimated posterior $p_{\bar{{\bm{\Psi}}}|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|\bm{\Theta}^{t},\bm{y})$ labelled ‘AMP’, and the limiting exact posterior $p_{\bar{{\bm{\Psi}}}|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},{\bm{\Psi}},{\bm{\varepsilon}}))$ labelled ‘Theory’, averaged over $30$ trials with true change point locations at $n/3$ and $8n/15$. The posterior is computed over the grid $[0,1]$ subsampled by a factor of ten. The experiment uses $p=800,{\mathbb{P}}_{\bar{\bm{B}}}={\sf N}({\bm{0}},\bm{I}),\sigma=0.1,\Delta=n/5,L=L^{*}=3$, with a uniform prior over all valid configurations. We observe that the estimated posterior closely matches the limiting posterior and concentrates around the ground truth as $\delta$ increases.
@@
Figures 3a and 3b compare the performance of AMP against four state-of-the-art algorithms: the dynamic programming ($\mathtt{DP}$) approach in (Rinaldo et al., 2021); dynamic programming with dynamic updates $(\mathtt{DPDU})$ (Xu et al., 2024); divide and conquer dynamic programming ($\mathtt{DCDP}$) (Li et al., 2023); and the complementary-sketching-based algorithm $\mathtt{charcoal}$ from Gao and Wang (2022). Hyperparameters are chosen using cross validation (CV), as outlined in Section A.1 of Li et al. (2023). The first three algorithms, designed for sparse signals, combine LASSO-type estimators with partitioning techniques based on dynamic programming. The $\mathtt{charcoal}$ algorithm is designed for the setting where the difference ${\bm{\beta}}^{(\eta_{\ell})}-{\bm{\beta}}^{(\eta_{\ell+1})}$ between adjacent signals is sparse. None of these algorithms uses a prior on the change point locations, unlike AMP which can flexibly incorporate both priors via ${\mathbb{P}}_{\bar{\bm{B}}}$ and $\pi_{\bar{{\bm{\Psi}}}}$. Figure 3a uses $p=200,\sigma=0.1,\Delta=n/10,L^{*}=L=3$ and a sparse Bernoulli-Gaussian signal prior ${\mathbb{P}}_{\bar{\bm{B}}}=0.5{\sf N}({\bm{0}},\delta\bm{I})+0.5\delta_{\bm{0}}$. AMP assumes no knowledge of the true sparsity level 0.5 and estimates the sparsity level using CV over a set of values not including the ground truth (details in Appendix E). Figure 3b uses $p=300,\Delta=n/10,L^{*}=L=3$ and a Gaussian sparse difference prior with sparsity level $0.5$ (described in (123)). AMP is run assuming a mismatched sparsity level of 0.9 and a mismatched magnitude for the sparse difference vector (details in Appendix E). Figures 3a and 3b show that AMP consistently achieves the lowest Hausdorff distance among all algorithms and outperforms most algorithms in runtime. Figures 8 and 9 in Appendix E show results from an additional set of experiments comparing AMP with $\mathtt{DCDP}$ (the fastest algorithm in Figure 3a). Figure 8 shows the performance of AMP with different change point priors $\pi_{\bar{{\bm{\Psi}}}}$ and suboptimal denoising functions such as soft thresholding. Figure 9 demonstrates the favourable runtime scaling of AMP over $\mathtt{DCDP}$ with respect to $p$, due to the LASSO computations involved in $\mathtt{DCDP}$.
@@
In Figure 4, we consider noiseless compressed sensing, where the signals $\{{\bm{\beta}}^{(i)}\}_{i\in[n]}$ are rotated versions of a $(255,255)$ sparse grayscale image used by Schniter and Rangan (2014). The fraction of nonzero components in the image is 8645/50625. We downsample the original image by a factor of three and flatten, yielding an operation dimension of $p=85^{2}=7225$. We set $\{{\bm{\beta}}^{(i)}\}_{i=1}^{0.3n}$ to be the image, $\{{\bm{\beta}}^{(i)}\}_{i=0.3n}^{0.7n}$ to be a $30^{\circ}$ rotated version, and $\{{\bm{\beta}}^{(i)}\}_{i=0.7n}^{n}$ to be a $45^{\circ}$ rotated version. We run AMP with a Bernoulli-Gaussian prior ${\mathbb{P}}_{\bar{\bm{B}}}$, $\Delta=n/4$ and $L=L^{*}=3$. Figure 4 shows image reconstructions along with the approximate posterior $p_{\bar{{\bm{\Psi}}}|\bar{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|\bm{\Theta}^{t},\bm{y})$. The approximate posterior concentrates around the true change point locations as $\delta$ increases, even when the image reconstructions are approximate. The experiment took one hour to complete on an Apple M1 Max chip, whereas competing algorithms did not return an output within 2.5 hours, due to the larger signal dimension compared to Figure 3a ($p=7225$ vs $p=200$).